# Configuration Requirements for 21-cm Forest Background Quasar Searches with the Moon-based Interferometer

### v1

@@ I Introduction
The 21-cm signal from the hyperfine transition of neutral hydrogen (HI) has become a powerful cosmological probe for the study of the early Universe, in particular, the Dark Ages, the Cosmic Dawn, and the Epoch of Reionization (EoR) [1]. Recently, Shao et al. [2] proposed a novel cosmological probe known as the 21-cm forest, which provides a unique probe of small-scale structures during the EoR. The 21-cm forest relies on the absorption characteristics of HI in the intervening structures along the line of sight with respect to the spectra of the bright background radio sources, such as radio-loud quasars. This technique provides valuable insights into the nature of dark matter (DM) and its role in shaping small-scale structure formation [3, 4, 5].
Breakthroughs in the 21-cm forest have been difficult to achieve because of ongoing challenges in both analytical modeling and observations. Because of the absence of analytical models, the parameter inference of the 21-cm forest relies on complex small-scale simulations. The substantial computational cost of simulations poses a significant challenge to constrain the astrophysical parameters. Recently, several new approaches have been proposed to address this challenge, including the deep-learning-driven likelihood-free parameter inference methods [6] and a halo-model-based analytical model for the one-dimensional power spectrum of the 21-cm forest [7].
Moreover, the detection of the 21-cm forest power spectrum remains challenging due to the limited sensitivity of current low-frequency observational facilities. This constraint is especially critical for the identification of high-redshift radio-loud quasars (HzRLQs), which serve as essential background sources. While optical and near-infrared surveys have extended quasar observations to redshifts beyond 6 [8, 9], the number of detected radio-loud quasars at such high redshifts remains small. A recent study [10] examined the abundance of HzRLQs and confirmed that the observations are still constrained by flux limitations even with the next-generation low-frequency radio telescope arrays, such as the low-frequency Square Kilometre Array (SKA-Low). This scarcity presents a significant obstacle to advancing 21-cm forest studies.
@@
The paper is organized as follows: In Section II, we describe the construction of the luminosity function for predicting the abundance of HzRLQs and the improvement of the model by including the obscuration effects. In Section III, we present various scenarios for the construction of the low-frequency radio array on the Moon. In Section IV, we present our results, including the influence of obscuration effects on the prediction of high-redshift quasars and how to build future moon-based arrays with the most significant detection of quasars and cost savings. Finally, in Section V, we summarize our conclusions and outline future research directions.
@@ II.1 Physical-Driven Model
    Calculate the black hole (BH) mass function. Quasars are driven by the active behavior of BHs at the centers of galaxies. Therefore, the first step is to determine the abundance of BHs in galaxy centers. We use a dark matter halo mass function in Sheth–Tormen form [20] to describe the abundance of halos of different masses at different redshifts. We assume that each halo hosts a galaxy and that each galaxy contains a central BH. The BH mass function is then derived from the halo mass function using the mass relationship between BHs ($M_{\rm BH}$) and their corresponding halos ($M_{\rm h}$) [21, 10],
@@
    Determine the quasar duty cycle with quasar luminosity function (QLF). The quasar duty cycle $D_{\rm q}=t_{\rm q}/t_{H(z)}$ quantifies the fraction of black holes in the quasar phase, where $t_{\rm q}$ is the quasar lifetime and $t_{H(z)}$ is the cosmic age at redshift $z$. Assuming Eddington-limited accretion, we constrain $t_{\rm q}$ by fitting observational data of the QLF at high redshifts [22, 23].
@@
    Determine the fraction of radio-loud quasars. We assume that $10\%$ of quasars are radio-loud [according to observational experience, e.g. 24, 25, 26] The fraction of radio-loud quasars over a flux threshold $F_{\rm th}$ is determined by integrating the radio-loudness distribution function $N(R)$,
@@ II.2 Obscuration-corrected QLF
where $\phi_{0}=0.73$, $\psi_{\rm max}=0.84$, $\psi_{\rm min}=0.2$, $\beta=0.24$, $L_{X,0}=10^{43.75}{\rm erg\,s}^{-1}$ [30, 31], The X-band luminosity is estimated via the bolometric correction factor of the hard X-ray band [32]:
@@ III Moon-based interferometer configuration
Literature research indicates that a significant detection of HzRLQs requires a radio interferometer with a sensitivity comparable to that of SKA-Low [10]. SKA-Low is designed to probe the early Universe through radio observations. It consists of 131 072131072131\,072131 072 log-periodic dipole antennas, grouped into $512$ stations, each containing $256$ antennas  [^1] https://www.skao.int/en/explore/telescopes/SKA-Low. Approximately $50\%$ of these stations are concentrated within a $1\,{\rm km}$ diameter central core, while the rest are distributed along three spiral arms extending up to $74\,{\rm km}$. The total effective collecting area is about 419 $419\,000\,{\rm m}^{2}$, providing exceptional sensitivity to faint radio signals. Operating within a frequency range of $50\,{\rm MHz}$ to $350\,{\rm MHz}$, SKA-Low can observe the 21-cm signal at redshifts approximately between $z\sim 3$ and $z\sim 27$, covering key cosmic epochs such as the EoR and the Cosmic Dawn.
@@
In this work, we consider the surveys are carried out with a continuum survey, which has the flux measurements integrated over the frequency band of $10\,{\rm MHz}$ to enhance the signal-to-noise ratio for continuum source detection. Due to the limited frequency resolution, the following 21-cm forest analysis requires additional deep-field observations. We also consider a 21-cm forest survey with spectrum resolution of $5\,{\rm kHz}$. Such high-frequency resolution significantly reduces the detection ability for HzRLQs, but it is fine enough for the 21-cm forest analysis without additional follow-up observations.
We assume $T_{\rm rx}=0.1T_{\rm sky}+40\rm K$ [33] and also assume a constant total survey area $S_{\rm tot}=10000\,{\rm deg^{2}}$ and total observation time $t_{\rm tot}=1\,{\rm yr}$. The observational noise variance may vary according to different effective collecting areas and survey efficiency.
@@
[…]
@@ IV.1 Configuration requirements with continuum survey
Figure 3: The number of HzRLQs detected with continuum survey as a function of redshift with different diameters of station and number of station configurations. The results with $\{(N_{\rm st},D_{\rm st})\}=\{(512,40\,{\rm m}),(2048,20\,{\rm m}),(8129,10\,{\rm m})\}$ are shown in different colors.
@@ IV.2 Configuration requirements with 21-cm forest survey
IV.3 Challenges in construction
@@ V Conclusion
We evaluate detection performance across a range of array sizes. For the smallest configuration considered ($N_{\mathrm{st}}=8$ stations with $40\,{\rm m}$ diameter), continuum surveys can detect HzRLQs up to $z\sim 10$, but the 21-cm forest survey becomes ineffective beyond $z\sim 5$. For SKA-scale arrays ($N_{\mathrm{st}}=512$), continuum surveys can detect quasars out to $z\sim 16$, while 21-cm forest observations remain feasible up to $z\sim 10$. For an extremely large array with $N_{\mathrm{st}}=2048$, the detection limit of the 21-cm forest survey extends to $z\sim 11$, supporting deeper exploration of the EoR.
[…]

### v2

@@ I Introduction
The 21-cm signal from the hyperfine transition of neutral hydrogen (HI) has become a powerful cosmological probe for the study of the early Universe, in particular, the Dark Ages, the Cosmic Dawn, and the Epoch of Reionization (EoR) [1]. The 21-cm absorption features imprinted by HI in the early Universe on the spectra of high-redshift radio-loud background sources has emerged as a promising probe of the intergalactic medium (IGM) during the Cosmic Dawn and EoR [2, 3, 4]. Recently, Shao et al. [5] proposed a novel cosmological probe known as the 21-cm forest, which provides a unique probe of small-scale structures during the EoR. The 21-cm forest relies on the absorption characteristics of HI in the intervening structures along the line of sight with respect to the spectra of the bright background radio sources, such as radio-loud quasars. This technique provides valuable insights into the nature of dark matter (DM) and its role in shaping small-scale structure formation [6, 7, 8, 9, 10].
Breakthroughs in the 21-cm forest have been difficult to achieve because of ongoing challenges in both analytical modeling and observations. Because of the absence of analytical models, the parameter inference of the 21-cm forest relies on complex small-scale simulations[11, 12]. The substantial computational cost of simulations poses a significant challenge to constrain the astrophysical parameters. Recently, several new approaches have been proposed to address this challenge, including the deep-learning-driven likelihood-free parameter inference methods [13] and a halo-model-based analytical model for the one-dimensional power spectrum of the 21-cm forest [14].
Moreover, the detection of the 21-cm forest power spectrum remains challenging due to the limited sensitivity of current low-frequency observational facilities. This constraint is especially critical for the identification of high-redshift radio-loud quasars (HzRLQs), which serve as essential background sources. While optical and near-infrared surveys have extended quasar observations to redshifts beyond 6 [15, 16, 17], the number of detected radio-loud quasars at such high redshifts remains small. A recent study [18] examined the abundance of HzRLQs and confirmed that the observations are still constrained by flux limitations even with the next-generation low-frequency radio telescope arrays, such as the low-frequency Square Kilometre Array (SKA-Low). This scarcity presents a significant obstacle to advancing 21-cm forest studies.
@@
The paper is organized as follows: In Section II, we describe the construction of the luminosity function for predicting the abundance of HzRLQs and the improvement of the model by including the obscuration effects. In Section III, we present various scenarios for the construction of the low-frequency radio array on the Moon. In Section IV, we present our results, including the influence of obscuration effects on the prediction of high-redshift quasars, how to build future moon-based arrays with the most significant detection of quasars and cost savings, and the baseline requirements to minimize the confusion effects. Finally, in Section V, we summarize our conclusions and outline future research directions.
@@ II.1 Physical-driven model
    Calculating the black hole (BH) mass function. Quasars are driven by the active behavior of BHs at the centers of galaxies. Therefore, the first step is to determine the abundance of BHs in galaxy centers. We use a dark matter halo mass function in Sheth–Tormen form [28] to describe the abundance of halos of different masses at different redshifts. We assume that each halo hosts a galaxy and that each galaxy contains a central BH. The BH mass function is then derived from the halo mass function using the mass relationship between BHs ($M_{\rm BH}$) and their corresponding halos ($M_{\rm h}$) [29, 18],
@@
    Determining the quasar duty cycle with quasar luminosity function (QLF). The quasar duty cycle $D_{\rm q}=t_{\rm q}/t_{H(z)}$ quantifies the fraction of black holes in the quasar phase, where $t_{\rm q}$ is the quasar lifetime and $t_{H(z)}$ is the cosmic age at redshift $z$. Assuming Eddington-limited accretion, we constrain $t_{\rm q}$ by fitting observational data of the QLF at high redshifts [30, 31].
@@
    Determining the fraction of radio-loud quasars whose radio flux exceeds the required threshold $F_{th}$. We assume that $10\%$ of quasars are radio-loud [according to observational experience, e.g. 32, 33, 34]. To account for the radio flux distribution of the quasar population at a given optical luminosity, we adopt the observed distribution of radio-loudness $R\equiv{\rm log_{10}}\left(L_{\rm 5GHz}/L_{4400}\right)$, where $L_{\rm 5GHz}$ and $L_{4400}$ are the luminosity measured at radio band ($5\,{\rm GHz}$) and optical band ($4400\,{\rm\AA }$), respectively. We use a radio-loudness distribution function $N(R)$ constrained using the observation data from Bariuan et al. [35]. The fraction of radio-loud quasars over a flux threshold $F_{\rm th}$ is determined by integrating the radio-loudness distribution function $N(R)$,
@@ II.2 Obscuration-corrected QLF
where $\phi_{0}=0.73$, $\psi_{\rm max}=0.84$, $\psi_{\rm min}=0.2$, $\beta=0.24$, $L_{\rm X,0}=10^{43.75}{\rm erg\,s}^{-1}$ [39, 40], The X-band luminosity is estimated via the bolometric correction factor of the hard X-ray band [41]:
@@ III Moon-based interferometer configuration
Literature research indicates that a significant detection of HzRLQs requires a radio interferometer with a sensitivity comparable to that of SKA-Low [18]. SKA-Low is designed to probe the early Universe through radio observations. It consists of $131072$ log-periodic dipole antennas, grouped into $512$ stations, each containing $256$ antennas  [^1] https://www.skao.int/en/explore/telescopes/SKA-Low. Approximately $50\%$ of these stations are concentrated within a $1\,{\rm km}$ diameter central core, while the rest are distributed along three spiral arms extending up to $74\,{\rm km}$. The total effective collecting area is about $419000\,{\rm m}^{2}$, providing exceptional sensitivity to faint radio signals. Operating within a frequency range of $50\,{\rm MHz}$ to $350\,{\rm MHz}$, SKA-Low can observe the 21-cm signal at redshifts approximately between $z\sim 3$ and $z\sim 27$, covering key cosmic epochs such as the EoR and the Cosmic Dawn.
@@
In this work, we consider that the surveys are carried out with a continuum survey, which has the flux measurements integrated over the frequency band of $10\,{\rm MHz}$ to enhance the signal-to-noise ratio for continuum source detection. An integration bandwidth of 10 MHz was selected to achieve an optimal balance between resolving the continuum’s spectral shape and maintaining sufficient detection sensitivity. This bandwidth is sufficiently narrow to capture spectral features while remaining broad enough to minimize noise. Due to the limited frequency resolution, the following 21-cm forest analysis requires additional deep-field observations. We also consider a 21-cm forest survey with spectrum resolution of $5\,{\rm kHz}$. Such high-frequency resolution significantly reduces the detection ability for HzRLQs, but it is fine enough for the 21-cm forest analysis without additional follow-up observations.
We assume $T_{\rm rx}=0.1T_{\rm sky}+40\,{\rm K}$ [42] and also assume a constant total survey area $S_{\rm tot}=10000\,{\rm deg^{2}}$ and total observation time $t_{\rm tot}=1\,{\rm yr}$. The observational noise variance may vary according to different effective collecting areas and survey efficiency. It is worth noticing that the sky temperature $T_{\rm sky}$ dominates the noise term at the low frequency bands, as shown in Equation (8). We systematically varied the receiver temperature to $T_{\rm rx}=0.1T_{\rm sky}+70\,{\rm K}$ and find that it is only $6.37\%$ more than the original. This analysis demonstrates that variations in receiver noise temperature have a negligible influence on the total system noise under these conditions.
@@
The solid angle of the field of view (FoV) for a single pointing is estimated using the primary beam size of the station given by $\Omega\sim(1.2\lambda/D_{\rm st})^{2}$, where $D_{\rm st}$ is the diameter of a station. For the subsequent analysis, we use the wavelength $\lambda$ corresponding to $150$ MHz to estimate the FoV. By varying the station diameter, the FoV for a single pointing changes, thereby affecting the survey efficiency, while keeping the total survey area and observation time constant. To maintain consistency with the current SKA-Low, we ensure the total effective collecting area remains at $\sim 419000\,{\rm m}^{2}$, and varying the number of stations and the diameter of the station simultaneously. In particular, we consider the following station configurations: $\{(N_{\rm st},D_{\rm st})\}=\{(512,40\,{\rm m}),(2048,20\,{\rm m}),(8192,10\,{\rm m})\}$.
@@ IV.1 Configuration requirements with continuum survey
Figure 3: The number of HzRLQs detected with continuum survey as a function of redshift with different diameters of station and number of station configurations. The results with $\{(N_{\rm st},D_{\rm st})\}=\{(512,40\,{\rm m}),(2048,20\,{\rm m}),(8192,10\,{\rm m})\}$ are shown in different colors.
@@ IV.2 Configuration requirements with 21-cm forest survey
IV.3 onfusion limit
Figure 6: Power-law fitting of source counts at $5.5\lesssim z\lesssim 20.5$ and flux density at $200\rm MHz$ based on SKA-Low.
Figure 7: The black line is the curve of the confusion limit $\sigma_{c}$ as a function of the baseline length $B$. Different color lines correspond to the observational noise under different configuration schemes.
The confusion limit is reached when the density of sources brighter than the observational noise becomes high enough within the area of the synthesized beam. At this point, the ability of the radio array to resolve individual sources is limited not by sensitivity, but by resolution. Consequently, the baseline length of the interferometric array plays a critical role in determining the array’s effectiveness in mitigating source confusion. With the luminosity function constructed in Section II, we simulate the radio source samples with flux density ranging from $10^{-4}\sim 10^{-1}{\rm Jy}$, which is large enough to cover the faint end of the radio sources, within the redshift of $5.5<z<20.5$, and estimate the column number density as a function of flux density. We fit the flux-density cumulative column number density with a power law function,
$N(>S)=CS^{\beta},$ (10)
where $N(>S)$ represents the column number density with sources’ flux density larger than $S$, and a power law function is characterized by the free parameters $C$ and $\beta$. The best-fit values are $C=3.94$, $\beta=-1.07$. The simulated flux-density cumulative column number density, as well as the best-fit curve, are shown in Figure 6.
The confusion limit is then estimated via [43, 44],
$\sigma_{c}=(\theta^{2}mC)^{-1/\beta},$ (11)
where $C=3.94$, $\beta=-1.07$ are the parameters of the best-fit power law function, $m=30$ is the threshold parameter [45], and $\theta$ is the angular resolution. The angular resolution of the interferometer array is related to the maximum length of the baseline, e.g. $\theta=c/(\nu B_{\rm max})$, where $c$ is the speed of light and $\nu$ is the corresponding frequency. To simplify the estimation, we ignore the frequency dependency and use $\nu=200\,{\rm MHz}$ in the following analysis.
We show the confusion limit as a function of the maximum baseline lengths in Figure 7. The noise level with different numbers of $40\,{\rm m}$ stations is shown with horizontal dashed lines in different colors. It is evident that increasing the maximum baseline length significantly reduces the confusion limit. To ensure that the confusion limit remains below the system’s thermal noise level, a minimum threshold for the maximum baseline should be established for configurations with varying numbers of stations. In general, a maximum baseline length of at least approximately $25\,{\rm km}$ is sufficient to render the confusion limit negligible.
IV.4 Challenges in construction
@@
Given that the data transfer rate of SKA-Low is as high as $7.15\,\rm Tbps$ [50] and the size of the data volume grows exponentially with the number of baselines ($N_{\rm st}(N_{\rm st}-1)/2$), such a large amount of data makes it technically challenging to transmit raw observation data from the Moon-based interferometer directly back to the Earth. To address the challenges of data transmission and real-time data processing in Moon exploration, we propose to establish a dedicated Moon-based data processing center. This facility will enable in situ data analysis, significantly reducing reliance on Earth-based processing while optimizing the science return from lunar missions.
@@ V Conclusion
We evaluate detection performance across a range of array sizes. For the smallest configuration considered ($N_{\mathrm{st}}=8$ stations with $40\,{\rm m}$ diameter), continuum surveys can detect HzRLQs up to $z\sim 10$, but the 21-cm forest survey becomes ineffective beyond $z\sim 5$. For SKA-scale arrays ($N_{\mathrm{st}}=512$), continuum surveys can detect quasars out to $z\sim 16$, while 21-cm forest observations remain feasible up to $z\sim 10$. For an extremely large array with $N_{\mathrm{st}}=2048$, the detection limit of the 21-cm forest survey extends to $z\sim 11$, supporting deeper exploration of the EoR. To minimize the confusion effects, the maximum baseline length needs to be $\gtrsim 25\,{\rm km}$.
We further explore configurations that maintain a fixed total collecting area ($\sim 419000\,\mathrm{m}^{2}$) but increase the number of stations by reducing the station diameter. This strategy improves field-of-view and survey efficiency, enabling higher detection rates without additional collecting areas. For example, replacing the baseline configuration ($N_{\mathrm{st}}=512$, $D_{\mathrm{st}}=40$ m) with $(N_{\mathrm{st}}=8192,D_{\mathrm{st}}=10\,\mathrm{m})$ substantially increases the number of detectable HzRLQs and raises the detection redshift limit. However, this improvement comes with greater demands on real-time data transmission, onboard storage, calibration, and computing infrastructure, particularly challenging for a Moon-based observatory.